# A dataset and model for recognition of audiologically relevant environments for hearing aids: AHEAD-DS and YAMNet+

### v3

@@ 4.1 Data Processing Workflow
For every five speech and environment sound file pairs, we mix them at 10,5,0,5,10-10,-5,0,5,10 dB signal-to-noise ratio (SNR) respectively. When speech and environment sounds were mixed, we boosted the levels of the quieter sound until it had equal RMS to the louder sound (this procedure occurs on top of standardisation). During mixing, speech was used as the signal and the environment sounds were used as noise. The speech levels were adjusted to the required SNR. The sounds of cocktail party and interfering speakers were used without mixing. No speech clip was used more than once.
The levels on the quieter sounds were boosted, as several sound clips for quiet indoors had very low RMS. When speech was mixed at 10-10 dB SNR, only near zero values were added. There were no or minimal signal changes after quantisation to 16-bit integers, when the mixed sound clips were saved back to disk. This causes sound clips with imperceptible speech being labelled with speech in quiet indoors, which is detrimental to model performance. The level boost ensures sound clips have acoustic realism and avoid labelling issues. Without the level boost, a human listener would not be able to determine there was speech in many speech in quiet indoors. With the level boost, background sounds are still nearly inaudible on manually listening, so the background acoustic signature is preserved.
@@
After standardisation and mixing, we divide sounds for each label into training, validation, and testing sets using a percentage split of 70, 10, and 20 respectively.
Figure 1: A data flow diagram showing the data processing procedures. Half of the environment sounds were mixed with speech.
4.2 Data Augmentation
Data augmentation was applied as a regularisation technique [29]. It was applied on-the-fly only during training. Three augmentation procedures were applied, each independently with a probability of fifty percent: a gain of between -6 to 6 dB; noise to perturb the data, from a uniform distribution within range -0.003 to 0.003; and the length of the clip is shrunk or stretched between 0.9 and 1.1, using the Fourier resampler [25]. These augmentation procedures were chosen because they were fast to implement, and parameter ranges were small to perturb the data, without changing it so much that on manual listening it is unrecognisable from the original. In a later section we find augmentation improves performance (See Label Smoothing, Augmentation, Unfreezing Only the Last Layer, and Random Weight Initialisation section).
5 YAMNet+ Sound Recognition Model
@@
The architecture diagram of YAMNet+ is shown in Figure 2. YAMNet+ is based on published YAMNet models [38] with the same internal architecture. We created a complete workflow for YAMNet+, as YAMNet was released with model definitions and weights but missing other components such as training code. We have released all code, making the implementation fully open source. For YAMNet+ we created training code, code to load weights, code for training via transfer learning, testing code, and conversion code to downsampling the model for deployment to edge computing platforms (We measured model speed using our test platform, a Google Pixel 3). YAMNet+ takes the waveform as input. A log mel spectrogram is computed from the waveform. The spectrogram is treated as a bitmap and processed using MobileNet [13]. The MobileNet starts with a convolutional block and a series of 13 blocks containing a depthwise separable convolution, batch normalisation, ReLU activation, convolution, batch normalisation and ReLU activation. There is a penultimate global average pooling layer. The final layer is a fully connected layer containing sigmoid activation of length equal to the number of class labels.
@@
YAMNet+ slices the waveform into windows of length 960ms with a 480 millisecond overlap, therefore, a 10 second clip will be split into 20 windows. Zero padding is added to ensure the input waveform is of sufficient length. Sound recognition is performed on every 960 millisecond window; scores were not combined for a waveform divided across multiple windows; each window is treated separately. We use 960 milliseconds as it is a hyperparameter which is fixed by the use of transfer learning; this is essential to get good performance (See Label Smoothing, Augmentation, Unfreezing Only the Last Layer, and Random Weight Initialisation section). If a longer window is required, it is possible to work around the fixed length by adding the results from multiple windows and applying a softmax mathematical operation.
YAMNet was originally trained using AudioSet containing 521 class labels. Both YAMNet and YAMNet+ contain 3,751,369 parameters when the model is initialised using 521 class labels. There are 3,743,831 parameters when the model is initialised using 14 class labels, as used in HEAR-DS and AHEAD-DS.
@@
Training occurred for a maximum of 100 epochs. An Adam optimiser [16] was used, with a learning rate of 0.00001. A decay-on-plateau learning rate scheduler was used. After 3 epochs, if performance on the validation set did not improve, the learning rate decays by multiplying by 0.5 after 6 epochs, if performance on the validation set did not improve, the training cycle was terminated early. The order of the training data was shuffled after every epoch.
Focal loss [20] with $\alpha=0.25$ and $\gamma=2$, was used. Focal loss down-weights easy and common classes, to better handle unbalanced datasets (AHEAD-DS is unbalanced).
A data loader was implemented to cycle through the data during training. It converts sound clips and labels into the format used during training. The sound clips are in WAV format and the labels are stored in a spreadsheet. Testing code was implemented to compute the confusion matrix, precision, recall, and mean average precision.
In order to optimise YAMNet+ for deployment on edge devices, the full-precision model was converted to a half-precision 16-bit floating-point model. This reduces the model size to 6.19 MiB. All subsequent testing was done using the 16-bit floating-point model.
Figure 2: The architecture of YAMNet+. The model computes the log mel spectrogram from the waveform. This is passed to a MobileNet which assigns confidence values to each label.
5.1 Label Smoothing
Label smoothing [22] with a smoothing factor of 0.1, was applied to the loss function. Label smoothing is a regularisation technique, which works by subtracting a small value from 1 for true classes, and adding a small value to 0 for incorrect classes. In a later section we show that label smoothing improves performance (See Label Smoothing, Augmentation, Unfreezing Only the Last Layer, and Random Weight Initialisation section).
5.2 Transfer Learning, Weight Initialisation, and Frozen Layers
YAMNet+ was initialised using transfer learning [41], where the weights were initially trained using AudioSet. Two ways of applying transfer learning were tested with YAMNet+. One method is to freeze all except the last layer, alternatively parameters in all layers are trainable. In a later section we will see having all layers trainable results in better performance (See Label Smoothing, Augmentation, Unfreezing Only the Last Layer, and Random Weight Initialisation section).
5.3 System Setup
YAMNet+ was developed using Python and TensorFlow. The development system was running Debian 12, FFmpeg 5.1.6, Python 3.11.2, TensorFlow 2.15.1 and 2.17.0, TensorFlow IO 0.37.1, and Keras 3. We make use of existing model definition code from YAMNet; this code is not fully compatible with newer versions of TensorFlow. We use the older TensorFlow to load and save a YAMNet model into a format which is compatible with newer TensorFlow. Then we use the newer TensorFlow to further tune this model.
TensorFlow Lite was used to convert the model to 16 bits to reduce the memory and computation requirements. The TensorFlow Lite model was deployed through an Android phone application [21]. The Android application was written using Flutter 3.24.5 and Dart 3.5.4.
The edge computing test system was a Google Pixel 3 phone running Android 13. The development system used a i7-1195G7 processor, 32 GiB of memory and a solid-state drive.
@@ 5.3 System Setup
The models were trained using only the central processing unit (CPU). It was not possible to get graphics processing unit (GPU) acceleration due to compatibility issues with TensorFlow; however, GPU use would reduce training time significantly.
6 Test Results
@@
A mean Average Precision (mAP) of 0.83 was achieved. This metric was calculated by averaging precision values across thresholds from 0.0 to 1.0 (in 0.1 increments) for each label, and then averaging those results. A threshold was applied to determine the label for each 960 millisecond window and compared against the ground truth. This gives the true positive, false positive, true negative, and false negative values. A higher mAP indicates better performance.
@@
The confusion matrix indicates an accuracy of 0.93 was achieved. The confusion matrix was produced by computing the confidence of each label. The highest scoring label was compared to the ground truth. Each number in each cell represents the number of 960 millisecond windows assigned to a label. A higher number along the diagonals indicates better results.
A mAP of 0.83 and an accuracy of 0.93 indicates the model demonstrates very promising performance in scene recognition. There were a few labels which were misidentified a notable number of times. The labels speech in music and music were often misidentified during pauses in speech and only music could be heard. The same issue occurs for speech in traffic and in traffic. The label speech in vehicle was sometimes misidentified as speech in wind turbulence because the sound of rushing air entering the vehicle can often be heard.
@@
Figure 3: YAMNet+ precision and recall curves. A mAP of 0.83 was achieved. A threshold was applied to determine the label for each 960 millisecond window and compared against the ground truth. A larger area under the curve indicates better performance.
@@
Figure 4: YAMNet+ confusion matrix. An accuracy of 0.93 was achieved. The highest scoring label was compared to the ground truth. Each number in each cell represents the number of 960ms windows assigned to a label. A higher number along the diagonals indicates better results.
6.1 Learning Rate Sensitivity Analysis
The learning rate vs several validation metrics are shown in Figure 5. The best validation loss, mAP, and accuracy was achieved with a learning rate of 0.00001.
Learning rates of 0.001, 0.0001, 0.00001, and 0.000001 were validated. The previously described decay-on-plateau learning rate scheduler was used. A decay of 0.5 was applied after 3 epochs where validation does not improve, early termination after 6 epochs where validation does not improve. Tests were performed using transfer learning by initialsing with weights trained on AudioSet.
The following results were observed: validation losses of 1.0, 0.11, 0.11, and 0.12; mAP scores of 0.34, 0.82, 0.85, and 0.85; and accuracy values of 0.13, 0.93, 0.95, and 0.94. These outcomes correspond sequentially to learning rates of 0.001, 0.0001, 0.00001, and 0.000001.
The learning rate of 0.00001 emerged as the most favorable choice as it achieved the best overall results. The learning rate of 0.001 achieved the worst results; the resulting ML model was likely not practically useful. This was caused by the relatively large learning rate causing parameters to overshoot optimal values.
(a) Learning rate vs validation loss. Lower values are better. The Y-axis has a scale break, this prevents the substantial validation loss of 1.0 (at a learning rate of 0.001) from visually dominating the graph.
(b) Learning rate vs validation mAP. Higher values are better.
(c) Learning rate vs validation accuracy. Higher values are better.
Figure 5: The learning rate vs several metrics. The best validation loss, mAP, and accuracy was achieved with a learning rate of 0.00001.
6.2 Label Smoothing, Augmentation, Unfreezing Only the Last Layer, and Random Weight Initialisation
@@
The results of validating several configurations, by adjusting hyperparameters are shown in Figure 6. The best overall result occurs under the default configuration with validation loss, mAP, and accuracy of 0.11, 0.85, and 0.95 respectively. The default configuration had label smoothing enabled, augmentation enabled, all layers were trainable, and weight initialisation used weights trained on AudioSet instead of random initialisation. The learning rate was set to 0.00001 for all experiments in this section.
When label smoothing was not applied the ML model produced a validation loss, mAP, and accuracy of 0.02, 0.82, and 0.95 respectively (compared to 0.11, 0.85, and 0.95 when label smoothing was enabled). The configuration with no label smoothing produced a low validation loss of 0.02. However, this result did not lead to better mAP or accuracy; this occurs when predictions are wrong but still relatively close to the correct results. The application of label smoothing yielded improved overall results.
When augmentation was not applied the ML model produced validation loss, mAP, and accuracy of 0.12, 0.84, and 0.94 respectively (compared to 0.11, 0.85, and 0.95 when augmentation was enabled). Augmentation yielded a very small impact on performance.
When only the last layer remained unfrozen the ML model produced a validation loss, mAP, and accuracy of 0.18, 0.87, and 0.80 respectively (compared to 0.11, 0.85, and 0.95 when all layers were trainable). This configuration produced the highest mAP of 0.87. However, a high mAP can still be achieved even when there are many false negatives in the presence of few false positives. The default configuration had better validation loss and accuracy results. Therefore, having all layers trainable yielded improved overall results.
When using random weight initialisation, the ML model produced a validation loss, mAP, and accuracy of 0.53, 0.61, and 0.15 respectively (compared to 0.11, 0.85, and 0.95 when using weights trained on AudioSet). The metrics indicate the ML model produced with random weight initialisation was likely not practically useful. Initialising the weights, trained on AudioSet, through transfer learning was essential to achieving a performant ML model.
The implications of the small size of AHEAD-DS can be observed in the analysis of results using random weight initialisation. Random weight initialisation produced a high loss for both training and validation, 0.53 and 0.53 respectively. This indicates underfitting. The learning curves show that fine tuning weights trained on AudioSet is necessary, when faced with the limited training data in AHEAD-DS.
The tradeoff of using transfer learning is that several hyperparameters become fixed. Changing a hyperparameter impacts all parameters dependent on the said hyperparameter. A new model will need to be initialised and re-trained from scratch using AudioSet, before transfer learning can be applied. This makes it resource intensive to change hyperparameters such as extending the window length beyond 960ms.
(a) Configurations vs validation loss. Lower values are better. The configuration with no label smoothing produced the lowest validation loss. However, this result did not lead to better mAP or accuracy. This occurs when predictions are wrong but still relatively close to the correct results.
(b) Configurations vs validation mAP. Higher values are better. The configurations where only the last layer remained unfrozen produced the highest mAP. However, a high mAP can still be achieved even when there are many false negatives in the presence of few false positives. This configuration did not produce the highest accuracy and had high validation loss.
(c) Configurations vs validation accuracy. Higher values are better. The equal best accuracy, after rounding, was produced under the default configuration.
Figure 6: Configurations vs several validation metrics. The best overall result was produced under the default configuration with label smoothing enabled, augmentation enabled, all layers trainable, and weight initialisation with weights trained on AudioSet instead of random initialisation.
6.3 Gain Sensitivity Analysis
Tests were performed at different input gains. The validation mAP and accuracy are shown in Figure 7.
We validated with a gain of -20 to +20 dB in 5 dB increments. We observed best validation mAP and accuracy results when testing with 0 dB gain.Deviations in sound levels above or below the model’s anticipated range, as determined by its training data, led to diminished performance. This indicates adjustment of the levels is necessary to achieve optimal performance when applying the model to new datasets.

### v4

@@ II-B AHEAD-DS Data Processing Workflow
Every five speech and environment sound file pairs are mixed at 10,5,0,5,10-10,-5,0,5,10 dB signal-to-noise ratio (SNR). The mixing function is shown in Algorithm 1. The reason to mix speech at 10-10 dB SNR was to allow for the possibility to detect speech, though not intelligibility, even in edge cases that would challenge a human with good hearing. When speech and environment sounds were mixed, the levels of the quieter sound were boosted until it had equal RMS to the louder sound, this procedure occurs on top of standardisation. During mixing, speech was used as the signal and the environment sounds were used as noise. The speech levels were adjusted to the required SNR. The sounds of cocktail party and interfering speakers were used without mixing. No speech clip was used more than once.
The reason to boost the quieter of speech or environment sound is to avoid the following situation. When speech was mixed at 10-10 dB SNR with quiet environment sounds such as quiet indoors, speech was imperceptible. This caused sound clips with imperceptible speech being labelled with a speech in environment label e.g. speech in quiet indoors, which is detrimental to model performance. The level boost ensures sound clips match the label.
@@
After standardisation and mixing, sounds are divided into training, validation, and testing sets using a percentage split of 70, 10, and 20 respectively.
Figure 1: A data flow diagram showing the data processing workflow.
Algorithm 1 Pseudocode to mix speech with environment sound.
1:function Mix($speech,env,snr$)
2:  $speech_{rms}\leftarrow\text{RMS}(speech)$
3:  $env_{rms}\leftarrow\text{RMS}(env)$
4:  $\triangleright$ Prevent division by zero
5:  if $speech_{rms}==0$ then
6:   $speech_{rms}\leftarrow 1e-6$
7:  end if
8:  if $env_{rms}==0$ then
9:   $env_{rms}\leftarrow 1e-6$
10:  end if
11:  $\triangleright$ Normalise amplitudes
12:  if $speech_{rms}>env_{rms}$ then
13:   $speech_{boosted}\leftarrow speech$
14:   $env_{boosted}\leftarrow env\times\frac{speech_{rms}}{env_{rms}}$
15:  else
16:   $speech_{boosted}\leftarrow speech\times\frac{env_{rms}}{speech_{rms}}$
17:   $env_{boosted}\leftarrow env$
18:  end if
19:  $\triangleright$ Calculate gain from SNR
20:  $gain_{factor}\leftarrow 10^{\left(\frac{snr}{20}\right)}$
21:  $\triangleright$ Apply gain and mix
22:  $speech_{gained}\leftarrow speech_{boosted}\times gain_{factor}$
23:  $mixed\leftarrow speech_{gained}+env_{boosted}$
24:  return mixedmixed
25:end function
II-C OpenYAMNet Model
@@
The architecture diagram of OpenYAMNet is shown in Figure 2. OpenYAMNet is based on YAMNet [27] with the same internal architecture. A complete workflow for OpenYAMNet was created, as the original YAMNet was released with model definitions and weights but missing other components such as training code. All code has been released, making the implementation fully open source. For OpenYAMNet the code consists of training code, code to load weights, code for training via transfer learning, testing code, and conversion code to downsampling the model for deployment to edge computing platforms.
OpenYAMNet takes the waveform as input. A log mel spectrogram is computed from the waveform. The spectrogram is treated as two dimension input and processed using MobileNet [14]. MobileNet starts with a convolutional block and a series of 13 blocks containing a depthwise separable convolution, batch normalisation, ReLU activation, convolution, batch normalisation and ReLU activation. There is a penultimate global average pooling layer. The final layer is a fully connected layer containing sigmoid activation of length equal to the number of class labels.
@@
OpenYAMNet slices the waveform into windows of length 960ms with a 480 millisecond overlap, therefore, a 10 second clip will be split into 20 windows. Zero padding is added to ensure the input waveform is of sufficient length. Scene recognition is performed on every 960 millisecond window and each window is treated separately. The 960 millisecond window length is a hyperparameter which is fixed by the use of transfer learning.
Figure 2: The architecture of OpenYAMNet.
II-D OpenYAMNet Training Workflow
@@
OpenYAMNet weights were initialised using transfer learning [41], taken from YAMNet trained on AudioSet. Training occurred for a maximum of 100 epochs. Focal loss [19] with $\alpha=0.25$ and $\gamma=2$ was used as the loss function. An Adam optimiser [17] was used, with an initial learning rate of 0.00001. A decay-on-plateau learning rate scheduler was used. After 3 epochs, if performance on the validation set did not improve, the learning rate decays by multiplying by 0.5. After 6 epochs, if performance on the validation set did not improve, the training cycle was terminated. The order of the training data was shuffled at the end of every epoch.
Label smoothing [21] with a smoothing factor of 0.1 was applied as a regularisation function. Data augmentation [29] was also applied as a regularisation technique. Three augmentation procedures were applied, each independently with a probability of fifty percent: a gain of between -6 to 6 dB, noise to perturb the data, from a uniform distribution within range -0.003 to 0.003, and the length of the clip is shrunk or stretched between 0.9 and 1.1, using the Fourier resampler [24].
The window length of 960 milliseconds and overlap of 480 milliseconds causes each 10 second audio clip in AHEAD-DS to be divided into 20 windows. The dataset contains 6,980 training, 1,001 validation, and 1,987 testing clips. A total of 139,600 windows are used for training, 20,020 for validation, and 39,740 for testing.
II-E Software Environment
@@
OpenYAMNet was developed using Python and TensorFlow. For technical details, links to the code are provided in the Data availability statement. The models were trained using only the central processing unit (CPU). It was not possible to get graphics processing unit (GPU) acceleration due to compatibility issues with TensorFlow, however, GPU use would reduce training time significantly.
In order to optimise OpenYAMNet for deployment on edge devices, the full-precision model was converted to a half-precision 16-bit floating-point model. This was done using TensorFlow Lite. The model was deployed through a phone application [20] written using Flutter. The edge computing test system was a Google Pixel 3 running Android.
III Results
This section outlines the results of testing OpenYAMNet using AHEAD-DS. The analysis includes the model’s performance on precision, recall, and accuracy. It also compares the impact of various hyperparameters, evaluates the processing time on an edge platform, and benchmarks the model against several other ML models.
III-A Precision, Recall, and Accuracy
@@
A mAP of 0.86 was achieved. All mAP metrics in this paper were calculated by averaging precision values across thresholds from 0.0 to 1.0, in 0.01 increments and then averaging those over all labels. This gives the true positive, false positive, true negative, and false negative values. A higher mAP indicates better performance.
@@
The confusion matrix indicates an accuracy of 0.93 was achieved. The confusion matrix was produced by computing the confidence of each label. The highest scoring label was compared to the ground truth. A higher number along the diagonals indicates better results.
(a) OpenYAMNet precision and recall curves for cocktail party, interfering speaker, and several environments.
(b) OpenYAMNet precision and recall curves for speech in several environments.
@@
Figure 3: OpenYAMNet precision and recall curves. A mAP of 0.86 was achieved on the test set. Each 960 millisecond window contributed a prediction outcome. A larger area under the curve indicates better performance.
@@
Figure 4: OpenYAMNet confusion matrix. An accuracy of 0.93 was achieved on the test set. Each 960ms window contributed a prediction outcome. A higher number along the diagonals indicates better results.
III-B Learning Rate Sensitivity Analysis
To evaluate learning rate sensitivity, four values were validated: 0.001, 0.0001, 0.00001, and 0.000001. For each learning rate, models were trained three times with different random seeds to obtain the mean and standard deviation (SD) for key metrics: loss, mAP, and accuracy on the validation set. The previously described decay-on-plateau learning rate scheduler was used, and model weights were initialised via transfer learning with weights pre-trained on the AudioSet dataset.
The results showed that learning rates of 0.0001, 0.00001, and 0.000001 all produced very similar performance, with a mean loss of 0.11, mean mAP ranging from 0.85 to 0.88, and a mean accuracy of 0.94. The low SD, which rounded to zero for several metrics, highlighted the consistency of these results. The learning rate of 0.001 performed poorly with a mean validation loss of 2.0, mean validation mAP of 0.5, and Mean Validation Accuracy of 0.14. The geometric mean of the three well performing learning rates, 0.00001 was chosen as the default learning rate.
TABLE III: The mean and standard deviations for loss, mAP, and accuracy over three iterations of each learning rate on the validation set.
Learning Rate Mean Validation Loss $\pm$ SD Mean Validation mAP $\pm$ SD Mean Validation Accuracy $\pm$ SD
0.001 2.00 $\pm$ 1.09 0.50 $\pm$ 0.27 0.14 $\pm$ 0.01
0.0001 0.11 $\pm$ 0.00 0.85 $\pm$ 0.01 0.94 $\pm$ 0.00
0.00001 0.11 $\pm$ 0.00 0.88 $\pm$ 0.00 0.94 $\pm$ 0.01
0.000001 0.11 $\pm$ 0.00 0.88 $\pm$ 0.00 0.94 $\pm$ 0.00
III-C Ablation Study
@@
The results of an ablation study validating several configurations are shown in Table IV. Each metric was calculated by running 10 iterations using bootstrapping, where 200 samples were chosen with replacement from the validation set. The result under the default configuration achieved a mean validation loss, mAP, and accuracy of 0.11, 0.88, and 0.95 respectively. The default configuration had label smoothing enabled, augmentation enabled, all layers were trainable, and weight initialisation used weights pre-trained on AudioSet. The learning rate was set to 0.00001 for all experiments in this section.
The regularising techniques of label smoothing and augmentation were validated. The model produced results which were very close to the models which did use these techniques, approximately within 0.01 mAP. Augmentation by applying a larger gain range of $\pm$24 dB, or use of weighted class training produced a slightly lower mAP, 0.85, but similar accuracy. The lowest validation loss of 0.01 was achieved when label smoothing was disabled.
When only the last layer was trainable, the ML model produced a mean validation loss, mAP, and accuracy of 0.18, 0.91, and 0.80 respectively. When using random weight initialisation, the ML model produced a validation loss, mAP, and accuracy of 0.53, 0.60, and 0.15 respectively. Using the TAU Urban Audio Visual Scenes 2021 dataset for pre-training produced similar results to random initialisation, a mean validation loss, mAP, and accuracy of 0.53, 0.62, and 0.15 respectively.
Normalising audio to a fixed root mean square (RMS) produced a noticeable drop in mAP and accuracy, 0.75 and 0.84 respectively. Similarly automatic gain control (AGC) produced a noticeable drop in mAP and accuracy, 0.77 and 0.85 respectively.
TABLE IV: Ablation study, the mean and standard deviations for loss, mAP, and accuracy for several configurations on the validation set over 10 iterations.
Configuration Mean Validation Loss $\pm$ SD Mean Validation mAP $\pm$ SD Mean Validation Accuracy $\pm$ SD
Default 0.11 $\pm$ 0.00 0.88 $\pm$ 0.01 0.95 $\pm$ 0.01
No Label Smoothing 0.01 $\pm$ 0.00 0.88 $\pm$ 0.01 0.95 $\pm$ 0.01
No Augmentation 0.11 $\pm$ 0.00 0.87 $\pm$ 0.01 0.94 $\pm$ 0.01
Unfreeze Only Last Layer 0.18 $\pm$ 0.01 0.91 $\pm$ 0.00 0.80 $\pm$ 0.01
Random Weight Initialisation 0.53 $\pm$ 0.01 0.60 $\pm$ 0.00 0.15 $\pm$ 0.02
TAU pre-training 0.53 $\pm$ 0.01 0.62 $\pm$ 0.00 0.15 $\pm$ 0.02
Augmentation $\pm$24 dB 0.11 $\pm$ 0.00 0.85 $\pm$ 0.01 0.94 $\pm$ 0.01
Weighted Classes 0.11 $\pm$ 0.00 0.85 $\pm$ 0.01 0.94 $\pm$ 0.01
Normalised RMS=0.2 0.17 $\pm$ 0.01 0.75 $\pm$ 0.01 0.84 $\pm$ 0.01
Block AGC RMS=0.2 0.17 $\pm$ 0.01 0.77 $\pm$ 0.01 0.85 $\pm$ 0.01
III-D 2-Fold Cross Validation
A two fold cross validation was implemented, the results are shown in Table V. Each fold was further divided so that ten percent of the samples within the fold were used as a hold-out dataset during training. This was necessary as the training loop expects the presence of a validation set. This created a final 45/5/50 train/validation/test split, where 45/5 belong to one fold for training and validation, and 50 belongs to the other fold for testing.
The results show very similar performance between 2-fold cross validation and the 70/10/20 data split. Both methods achieved a loss of 0.11, mAP of 0.86, and accuracy of 0.93.
TABLE V: The loss, mAP, and accuracy for 2-fold cross validation using AHEAD-DS.
Fold mAP Accuracy
Fold 1 train/Fold 2 test 0.86 0.93
Fold 2 train/Fold 1 test 0.86 0.93
Mean 0.86 0.93
III-E Gain Sensitivity Analysis
Tests were performed at different input gains. The validation mAP and accuracy are shown in Figure 5. Validation was performed with a gain of -20 to +20 dB in 5 dB increments. Gain was applied to both speech and background noise, so SNR did not change.Close performance was observed in validation mAP and accuracy when testing with -5, 0, and 5 dB gain. A higher mAP was achieved at 0 dB due to fewer false positives, but the accuracy was lower compared to -5 and 5 dB due to more false negatives.